A plane inclined at angle theta has a velocity of 10:1. Calculate the inclination of the plane to the horizontal
1
Expert's answer
2020-11-30T14:52:51-0500
"\sin{\theta}=\frac{1}{10}\\\theta=\arcsin{\frac{1}{10}}=5.7\degree"